Calculate the polarizing angle for light traveling from water, of refractive index 1.33 to glass of refractive index 1.53


1
Expert's answer
2021-03-08T08:33:40-0500

Given,

Refractive index of water "(\mu_w) = 1.33"

Refractive index of glass "(\mu_g) = 1.53"

We know that reflected light is always polarized, and it reflects at "90^\circ"

"\tan\theta = \frac{\mu_2}{\mu_1}"

Now substituting the values,

"\Rightarrow \tan\theta = \frac{1.53}{1.33}"


"\Rightarrow \tan\theta = 1.15"

"\theta = \tan^{-1}(1.15)"

"=48.99^\circ"